OpenStudy (anonymous):

Let y = ln(x + 2) - ln5 find the derivative

OpenStudy (campbell_st):

the question can be simplifed to ln((x+2)/5) f(x) = (x+2)/5 f'(x) = 1/5 dy/dx = 1/5/(x+2)/5 = 1/(x+2)